OpenStudy (anonymous):

WILL GIVE MEDAL Given two points, write the equation of the line in slope-intercept form (-3,4),(1,-2) PLEASE SHOW WORK

OpenStudy (anonymous):

\[\frac{ -2-4 }{ 1-(-3) }=\frac{ -3 }{ 2 }\]

OpenStudy (anonymous):

This is your slope

OpenStudy (anonymous):

Plug in a set of coordinates into point intercept form y-4=-3/2(x-(-3))

OpenStudy (anonymous):

\[y=\frac{ -3 }{ 2 }x-\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

\[m=\frac{ -2-4 }{ 1-(-3) }=\frac{ -3 }{ 2 } \]

OpenStudy (anonymous):

\[(y _{1}-y _{2})=m(x _{1}-x _{2}) slope-intercept form\]

OpenStudy (anonymous):

\[\frac{ y^2-y^1 }{ x^2-x^1 }\] So: \[\frac{ -2-4 }{ 1-(-3) }= \frac{ -6 }{ 4 }=\frac{ -3 }{ 2 }\] Plug in -3/2 for m in the equation: y=mx+b y=-3/2x+b Then pick a point and plug in the x and the y: \[y=\frac{ -3 }{ 2 }x+b\] I'll use (1, -2): \[-2=\frac{ -3 }{ 2 }(1)+b\] Solve it out: \[-2=\frac{ -3 }{ 2 }+b\] Add -3/2 to -2 and your answer is: \[y=\frac{ -3 }{ 2 }x - \frac{ 1 }{ 2 }\]